OpenStudy (anonymous):

In parallelogram DEFG, the difference between two adjacent angles is 32 degrees. Which is the measure of the larger angle?

OpenStudy (anonymous):

x + y =180 x - y = 32 adding 2x = 212 x = 106

OpenStudy (anonymous):

larger angles = 106 degrees

OpenStudy (anonymous):

x+y=180 (adjacent angles in a parallelogram are supplementary) x-y=32 (given) if you subtract the equations, you get 2y=148, or y=74. x+y=180 x+74=180 x=106

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yep 74 = measure of smaller angle

OpenStudy (anonymous):

ok - plz click on good answer(s) as you see fit, so this will close the question